OpenStudy (anonymous):

given x=5, y=3/5, z=3.3 evaluate each expression 3^y

OpenStudy (thesecret20111):

That would be 3^3/5. Input this to your calculator.

OpenStudy (jhannybean):

\[\huge 3^{3/5} = \sqrt[5]{3^3}\]

OpenStudy (jhannybean):

You see how that works?

OpenStudy (anonymous):

A. 101 B.0.203 C.32.6 D.1.1

OpenStudy (jhannybean):

http://www.sketchtoy.com/37817495

OpenStudy (jhannybean):

now input \(\large \sqrt[5]{3^3}\) into your calculator and solve.

OpenStudy (anonymous):

it doesnt let me